# Topological protection breakdown: a route to frustrated ferroelectricity

### v2

@@ Abstract
Phases manifesting topological patterns in functional systems, like ferroelectric and ferromagnetic vortex superlattices, are typically frustrated non-ergodic states characterized by high-dimensional energy landscapes with intricate and apparently ungovernable behavior. This is also the case for potassium-tantalate-niobate (KTN) crystals. These transparent ferroelectrics manifest remarkable but little-understood metastable domain patterns at optical (micrometer and above) scales near the cubic-to-tetragonal structural phase transition. Here, we formulate the Topological Breakdown Model based on the competition between intrinsic scales of domain-domain collinear and non-collinear interactions associated with polarization-charge screening. The model is able to explain the KTN mesoscopic domain patterns and phase diagram as a function of temperature and external electric field. Findings include a precise set of sharp and broad percolative transitions that are experimentally verified, validating the model. Our study identifies the central role played by competing topologically protected states, identifying the fundamental link between topological protection and frustration that supports a hitherto unexplored functional non-ergodic arena.
A knot in a string is a remarkably stable physical form of information storage since it can only be erased by untangling it [1, 2, 3, 4]. Knot-like structures emerge naturally in the form of topological defects in a wide variety of materials [5], such as liquid-crystals [6], fluids [7], ferromagnets [8, 9], and ferroelectrics [10, 11]. The property of being resilient to small external perturbation on consequence of a predefined geometrical constraint is called topological protection [12]. Protected states are ubiquitous in condensed matter systems in the form of topological defects, such as skyrmions in ferromagnetic materials [13], and find applications in electromechanical and spintronic devices for information storage [14] and logic devices [15]. The same phenomenon is responsible for the stability of the intricate low-dimensional topological polar structures (e.g., quadrant domains [16], polar flux closures [17, 18], and vortices [19]) found in ferroelectric materials, that, in recent years, have garnered significant attention because of their stability [20, 21, 22].
At present, their potential as noise-resistant memory and processing elements is largely unexplored, especially because their behavior appears intimately connected with an intriguing but little-understood frustration in the hosting material [23, 12, 24]. In polar ferroelectrics, topological defects form spontaneously during the discrete inversion-symmetry-breaking phase-transition as closed-flux patterns [25, 18, 17], while the dielectric response is relaxor-like and non-ergodic [26, 27, 28, 29]. Typically, topological defects spontaneously emerge in proximity of a structural transition, where functional materials also manifest history-dependence [28]. Non-ergodicity plays a major role in many natural information-driven phenomena, such as memory effects [30], protein-folding mechanisms [31], brain dynamics [32, 33], and slow relaxation dynamics [34]. However, conventional explanations for non-ergodicity, including Griffiths phases and glassy states, fail to fully account for key experimental features in disordered ferroelectrics [35, 26]. In functional materials, in turn, non-ergodic behavior has never been fully understood and modelled, greatly curtailing potential applications. Specifically, understanding the behavior of topological structures, i.e., polarization vortices, and controllable domain patterns in ferroelectric and ferroic materials is key to developing innovative memory devices and processors, or even entirely new classes of neural mass-models and richer scenarios to enact machine-learning algorithms, optimally balancing accuracy and efficiency.
@@
We here formulate the Topological Breakdown Model (TBM) with the aim of finding the physical underpinnings of complex mesoscopic (from several micrometer to hundreds of micrometer) domain pattern behavior in ferroelectric KTN:Li below its room-temperature Curie point. The TBM is based on the competition between intrinsic mesoscopic scales through collinear and non-collinear interactions. These emerge geometrically out of a topologically protected fabric from discrete-inversion-symmetry-breaking and volume charge-screening flux-closure constraints [25, 36]. Computational analysis, real-space renormalization group (RG) techniques, and experimental results allow the understanding of how ensembles of flux-closure patterns and superlattices react to external stimuli. Predictions of the domain pattern phase diagram versus temperature and versus electric bias field identify a signature metastable state that is experimentally confirmed using percolation analysis. Our idea links directly topologically protected defects to frustration, demonstrating how competition between ordered topologically protected states generates an entirely new functional arena whose macroscopic result is the previously unexplained non-ergodicity.
I The Topological Breakdown Model (TBM) for Ferroelectrics
@@
We develop the TBM starting from the basic fact that the elementary cell of ordinary ferroelectric perovskites $(ABO_{3})$ above $T_{C}$ exhibits a cubic symmetry. In the vicinity of $T_{C}$, the B-ion features an atomic displacement along the four-fold axis, leading to a discrete inversion-symmetry breaking, which induces a distorted electron distribution and spontaneous electric polarization [37]. Phenomenologically, below $T_{C}$, a finite sample relaxes into a pattern of different ferroelectric domains, mesoscopic regions in which the polarization is parallel to a given crystal axis. While each domain enucleates independently with a specific macroscopic polarization $\mathbf{p}$, as these grow, regions with different spontaneous polarization meet, fusing together or forming topological defects in the form of domain walls. Based on Maxwell’s electrostatic equations, domains arrange to allow the screening of volume polarization charge, i.e., the minimization of $\nabla\cdot\mathbf{p}$ [38]. This leads to two types of defects (see Fig.1(a)), type i) walls which separate two domains with anti-parallel polarization with $\mathbf{p}_{i}=-\mathbf{p}_{j}$ and type ii) walls that form between orthogonal polarizations $\mathbf{p}_{i}\cdot\mathbf{p}_{j}=0$. For a given starting close-packed multi-domain distribution, dynamics appear driven by domain dipole-dipole interaction with the constraint that $\nabla\cdot\mathbf{p}\simeq 0$ on the walls. While dipole-dipole-like interaction, dominant across type i) defects, can be described through the paradigmatic Heisenberg model $(\mathcal{H}_{H}\propto J_{ij}\mathbf{p}_{i}\cdot\mathbf{p}_{j})$, the interaction across type ii) walls is governed by polarization charge screening, i.e., by non-collinear interactions.
@@
$\mathcal{H}=-\underset{\langle i,j\rangle}{\sum}J_{ij}\mathbf{p}_{i}\cdot\mathbf{p}_{j}+\underset{\langle i,j\rangle}{\sum}\mathbf{D}_{ij}\cdot(\mathbf{p}_{i}\times\mathbf{p}_{j})-\sum_{i}\mathbf{h}\cdot\mathbf{p}_{i}$ (1)
where $\mathbf{p}_{i}$ is the n-component spontaneous polarization distribution of the domain in the ”ith” site with discrete values $\mathbf{p}_{i}=(\pm p_{x},\pm p_{y})$, $J_{ij}=JW_{ij}$ are the elements of the $N\times N$ connectivity matrix W, with $J$ the global coupling strength, and $\mathbf{D}_{ij}\propto\mathbf{p}\times\mathbf{e}_{ij}$ where $\mathbf{e}_{ij}$ is the unit vector pointing from i-th site to j-th site and $\mathbf{p}$ the polarization vector [47]. The last term corresponds to an applied external field and is described, as usual, as $\mathcal{H}_{e}=-\sum_{i}\mathbf{h}\cdot\mathbf{p}_{i}$.
A first order parameter is taken to be the modulus $\left|\textbf{p}\right|_{i}$ of the polarization in each direction $P=\frac{1}{N}\sum_{i}\left|\textbf{p}\right|_{i}$, where $i$ represents the sum over nodes, and $|p|=\sqrt{p_{x}^{2}+p_{y}^{2}}$. A second order parameter is taken to be $v=\frac{1}{2N}\underset{\langle i,j\rangle}{\sum}\frac{|\mathbf{p}_{i}\times\mathbf{p}_{j}|}{|\mathbf{p}_{i}||\mathbf{p}_{j}|}$, this quantifying emergent rotational order in the system (i.e., vortex-like structures). We emphasize that our second proposal is an angular order parameter that captures the angular ordering of the system. Note that, unlike low-dimensional systems undergoing a continuous symmetry breaking, where the emergence of topological defects in the form of vortices is expected as a direct consequence of the Mermin-Wagner theorem [50, 51], for the present discrete symmetry case, vortex-like patterns are a direct consequence of the DMI.
@@ II.1 SC-tiling Intrinsic Scales
Figure 1(c) shows the so-called specific heat ($C$, see Appendix A) of the SC structure. Along with the plateau that reflects the 2D nature of the lattice, as $C=\frac{d}{2}$ for scale-invariant architectures [58], $C$ presents two possible small-scale modes or ultraviolet cut-offs (in the jargon of the Renormalization Group), $\Lambda_{1}$ and $\Lambda_{2}$, independent of the system size, plus a third peak at larger values of $\tau$ reflecting the whole lattice scale. Compared to a regular 2D lattice that only presents two peaks, here, the extra peak plays, as discussed below, a key role in determining, altering, and disrupting the phases and phase transitions of the system.
Figure 2: Topological symmetry-breaking. (a) Temperature phase diagram of Eq.(1) computed using direct simulations on a SC lattice with $N=1024$ nodes ($J=1$). Polarized states are computationally detected with the polarization order parameter, $P$, while vortex phases are detected through the vortex order parameter $v$. The metastable phase corresponds to the broad region of divergent polarization susceptibility $\chi_{P}=N\sigma^{2}_{P}$. (b)-(c) Polarization ($P$) and vortex ($v$) order parameters versus temperature for different lattice sizes (see legend, $N=4L^{2}$) for (b) $J/D=1.4$ and (c) $J/D=0.8$. Inset: Rescaled temporal variance $\chi_{P}=N\sigma^{2}_{P}$ versus temperature, $T$ for different system sizes. The metastable phase also presents characteristic temporal oscillations in the global system polarization (see Appendix B). (d) Top: Far-field imaging of a SC structure at $T$= $T_{C}$-2K. Bottom: Computational 2D Fourier transform of an ordered vortex phase ($J/D=0.8,~{}T=0.4,~{}N=1024)$. (e) Temperature phase diagram of the model versus electric field $h$ for $J/D=0.8$ and $N=1024$ nodes. The breakup of the vortex phase translates into the emergence of a metastable phase where the ferroelectric and vortex domains naturally coexist. (f)-(g) Polarization ($P$) and vortex ($v$) order parameters as a function of the electric field for different lattice sizes (as in (b) and (c)) for (f) $T=0.4$ and (g) $T=0.7$. The peak in the system susceptibility is indicated here as a vertical dashed line. Note how the critical field at low temperatures splits into two different phase transitions when T increases, with naturally coexisting vortices and ferroelectric domains. All curves have been averaged over $10^{2}-10^{3}$ independent realizations.
@@ II.2 Phase Diagram
Figure 2(a) shows the temperature phase diagram of the system as a function of the ratio between the dipole-dipole alignment interaction, $J$, and the DMI, $D$. Results are achieved through extensive Monte Carlo simulations of the TBM, running Eq.(1) on top of the SC-tiling. In order to reach zero local charge density, the constraint $\nabla\cdot\mathbf{p}_{\Lambda_{1}}=0$ is imposed as the energy is minimized (see Fig. 1(b)), centering it on each ”plaquette” (the elementary protected cell $\Lambda_{1}$) to avoid nonphysical situations (see Appendix B for case examples of microscopic configurations). Inspection of the phase diagram reveals two stable collective dynamical regimes: vortices and ferroelectric phases. In particular, the microscopic scale $\Lambda_{1}$ now supports the existence of an ordered domain mosaic of 2D protected vortex cores. Different types of bifurcation lines separate the phases of the system. In particular, for low temperatures, increasing the ratio $J/D$, a first-order phase transition separates the vortex phase from the ferroelectric one. Conversely, a second-order phase transition separates the vortex phase from the paraelectric one. Interestingly, a new phase emerges between the ferroelectric and paraelectric phases: a metastable phase where transient vortices prevent the system from reaching an ordered state, sequentially changing the full system orientation (see SM [55]).
@@
Figures 2(b) and (c) show order parameters as a function of the temperature for the polarization, $P$, and vortex, $v$, for two different values of the ratio $J/D$. Note that $P$ and $v$ correctly detect the stable ordered phases. Insets show the anomalous temporal fluctuations $\sigma^{2}_{P}$ and $\sigma^{2}_{v}$, which exhibit a pronounced peak located at the ($N$-dependent) transition point between the different phases. In particular, the broad divergence of $\sigma^{2}_{P}$ proves the emergence of the intermediate phase between the ferroelectric and the paraelectric ones (see also Appendix B).
@@
Figure 3: Direct imaging of SC ferroelectric clusters. Binarized crossed-polarizer transmission microscopy images for (a) $T_{C}$-4.5K and increasing electric field (0 kV/cm, 2.86 kV/cm, and 2.91 kV/cm, respectively, using $\theta=25$). Light transmission shows an abrupt transition at $E_{C}\approx$2.9 kV/cm. (b) Various conditions of temperature and bias electric field to enhance the critical point for the two first temperatures and the metastable region for $T_{C}$-2.5K (see images, with $\theta=37$, $37$, and $35$, respectively). (c-e) Light intensity versus applied electric field ($E$) for three different temperatures and different threshold values ($\theta$, see legend): (c) $T_{C}-4.5K$. An abrupt phase transition is present at a field $E_{C}\approx 2.9$ kV/cm for a wide range of $\theta$ values. (d) $T_{C}-3.5K$. The abrupt transition changes its nature to a continuous one with $E_{C}\approx$ 2.75 kV/cm and (e) $T_{C}-2.5K$. The system exhibits two different phase transitions at different critical fields, $E_{C}\approx$2.5 kV/cm and $E_{C}\approx$ 2.86 kV/cm (see also Appendix E).
@@
Findings are compared to experiments in SCs observed in KTN:Li crystals (see Appendices C and D). This material, which led to the first observation of wave replica-symmetry-breaking [59], has been developed by several groups [60, 61, 62, 63] as a lead-free optically transparent alternative to other commonly used perovskites [64]. Moreover, compared to commercially available crystals, such as BTO [53] and SBN [65], it has tunable properties with a room-temperature ferroelectric transition [66, 67, 68]. Fourier computational analysis of the vortex phase (see also SM [55]) is found to match observed optical far-field results (see Fig.2(d)), in agreement with the previous SC models based on an interlocked regular structure of spontaneous polarization [28, 69, 67]. The TBM model then provides a feasible explanation to why the SC phase appears only in some specific materials, as the internal structure must be fine-tuned to present a specific $J/D$ ratio.

### v4

@@ Abstract
Phases manifesting topological patterns in functional systems, like ferroelectric and ferromagnetic vortex superlattices, can manifest intricate and apparently ungovernable behavior, typical of frustrated non-ergodic states with high-dimensional energy landscapes. This is also the case for potassium-tantalate-niobate (KTN) crystals. These transparent ferroelectrics manifest remarkable but little-understood metastable domain patterns at optical (micrometer and above) scales near the cubic-to-tetragonal structural phase transition. Here, we formulate the Topological Breakdown Model based on the competition between intrinsic scales of domain-domain collinear and non-collinear interactions associated with polarization-charge screening. The model is able to explain observed KTN mesoscopic domain patterns and phase diagram as a function of temperature and external electric field. Findings include a precise set of sharp and broad percolative transitions that are experimentally verified, validating our model. Our study identifies the central role played by competing topologically protected states, identifying a fundamental link between topological protection and frustration that supports a hitherto unexplored functional non-ergodic arena.
Disordered ferroelectric crystals can manifest history-dependence and frustrated behavior in proximity of their structural phase transitions, features that affect response to various degrees, becoming dominant in so-called relaxor ferroelectrics [1, 2, 3, 4, 5, 6]. The regime forms a modelling challenge, as conventional frustration mechanisms, including Griffiths phases and glassy states, fail to fully account for key experimental features [3, 7]. In a subset of disordered ferroelectrics, including solid-solution potassium-tantalate-niobate (KTN), the complex relaxor-like behavior [3, 5] coincides with the emergence of domain ordering in the form of ferroelectric superlattices or supercrystals [8]. These patterns appear connected to built-in striations [9, 10] or engineered heterostructures [11, 12], coupled to the reduction of volume polarization charge $\nabla\cdot\mathbf{p}\simeq 0$, where $\mathbf{p}$ is the spontaneous polarization field. The constraint introduces closed-loop or flux-closure conditions, so that energetically preferred structures appear in the form of ferroelectric topological defects [13, 14]. These have topological protection, in that each domain in the pattern can be reoriented only at the expense of modifying all the others [15]. The question naturally arises if there is a connection between these topological features and the emergence of complexity-driven frustrated behavior. In fact, the volume charge screening condition makes $\mathbf{p}$ topologically analogous to a magnetic induction field, suggesting that domain-domain interaction can also be governed, at the mesoscale, by effective interaction terms that are typical of magnetic interaction. At present, no effective domain interaction model has been proposed to take into account of both standard collinear dipole-dipole interaction and effective magnetic-like noncollinear interaction, leaving the connection between closed-flux topologically-protected domain patterns and frustration largely unexplored.
@@
We here formulate the Topological Breakdown Model (TBM) with the aim of identifying the role of effective non-collinear interaction in determining complex mesoscopic (from several micrometer to hundreds of micrometer) domain pattern behavior in systems supporting ferroelectric superlattices. The model is based on the competition between intrinsic mesoscopic scales through both collinear and non-collinear interactions that emerge geometrically from discrete-inversion-symmetry-breaking and volume charge-screening flux-closure constraints [16, 17]. Computational analyses and real-space renormalization group (RG) techniques allow us to describe how ensembles of flux-closure patterns and superlattices react to external stimuli. Predictions of the domain pattern phase diagram versus temperature and versus electric bias field identify a frustrated metastable phase that is experimentally confirmed in near-transition KTN:Li using percolation analyses. Our findings link directly topologically protected defects to frustration and expand our ability to design and control ferroelectric response.
I The Topological Protection Breakdown Model (TBM) for Ferroelectrics
@@
We develop the TBM starting from the basic fact that the elementary cell of ordinary ferroelectric perovskites $(ABO_{3})$ above $T_{C}$ exhibits a cubic symmetry. In the vicinity of $T_{C}$, the B-ion features an atomic displacement along the fourfold axis, leading to a discrete inversion-symmetry breaking, which induces a distorted electron distribution and spontaneous electric polarization [18]. We note that this picture holds for many disordered perovskite ferroelectrics, such as KTN [19], but may not apply to the class of relaxor ferroelectrics, such as PMN, where well-define structural transitions are not observed [20, 21]. Phenomenologically, below $T_{C}$, a finite sample relaxes into a pattern of different ferroelectric domains, mesoscopic regions in which the polarization is parallel to a given crystal axis. While each domain enucleates independently with a specific macroscopic polarization $\mathbf{p}$, as these grow, regions with different spontaneous polarization meet, fusing together or forming topological defects in the form of domain walls. Based on Maxwell’s electrostatic equations, domains arrange to allow the screening of volume polarization charge, i.e., the minimization of $\nabla\cdot\mathbf{p}$ [22]. This leads to two types of defects (see Fig.1(a)), type i) walls which separate two domains with anti-parallel polarization with $\mathbf{p}_{i}=-\mathbf{p}_{j}$ and type ii) walls that form between orthogonal polarizations $\mathbf{p}_{i}\cdot\mathbf{p}_{j}=0$. For a given starting close-packed multi-domain distribution, dynamics appear driven by domain dipole-dipole interaction with the constraint that $\nabla\cdot\mathbf{p}\simeq 0$ on the walls. While dipole-dipole-like interaction, dominant across type i) defects, can be described through the paradigmatic Heisenberg model $(\mathcal{H}_{H}\propto J_{ij}\mathbf{p}_{i}\cdot\mathbf{p}_{j})$, the interaction across type ii) walls is governed by polarization charge screening, i.e., by non-collinear interactions.
@@
We want to emphasize that the term ”breakdown” refers here specifically to the loss of topological protection associated with flux-closure configurations. These configurations, stabilized by the local constraint $\nabla\cdot\mathbf{p}=0$, are dynamically robust and can only be destroyed via coordinated, nonlocal domain rearrangements, thus functionally mimicking topological defects in low-dimensional systems. The breakdown occurs when thermal fluctuations or applied external fields induce a transition from vortex-protected states to states with extended or reoriented domains, inducing the collapse of this protection mechanism.
@@
$\mathcal{H}=-\underset{\langle i,j\rangle}{\sum}J_{ij}\mathbf{p}_{i}\cdot\mathbf{p}_{j}+\underset{\langle i,j\rangle}{\sum}\mathbf{D}_{ij}\cdot(\mathbf{p}_{i}\times\mathbf{p}_{j})-\sum_{i}\mathbf{E}\cdot\mathbf{p}_{i}$ (1)
where $\mathbf{p}_{i}$ is the n-component spontaneous polarization distribution of the domain in the ”ith” site with discrete values $\mathbf{p}_{i}=(\pm p_{x},\pm p_{y})$, $J_{ij}=JA_{ij}$ are the elements of the $N\times N$ adjacency matrix $\hat{A}$, with $J$ the global coupling strength, and $\mathbf{D}_{ij}\propto\mathbf{p}\times\mathbf{e}_{ij}$ where $\mathbf{e}_{ij}$ is the unit vector pointing from i-th site to j-th site and $\mathbf{p}$ the polarization vector [31]. The last term corresponds to an applied external field and is described, as usual, as $\mathcal{H}_{e}=-\sum_{i}\mathbf{E}\cdot\mathbf{p}_{i}$.
A first order parameter is taken to be the modulus $\left|\textbf{p}\right|_{i}$ of the polarization in each direction $P=\frac{1}{N}\sum_{i}\left|\textbf{p}\right|_{i}$, where $i$ represents the sum over nodes, and $|p|=\sqrt{p_{x}^{2}+p_{y}^{2}}$. A second order parameter is taken to be $v=\frac{1}{2N}\underset{\langle i,j\rangle}{\sum}\frac{|\mathbf{p}_{i}\times\mathbf{p}_{j}|}{|\mathbf{p}_{i}||\mathbf{p}_{j}|}$, this quantifying emergent rotational order in the system (i.e., vortex-like structures). This is an angular order parameter that captures the angular ordering of the system. Note that, unlike low-dimensional systems undergoing a continuous symmetry breaking, where the emergence of topological defects in the form of vortices is expected as a direct consequence of the Mermin-Wagner theorem [34, 35], for the present discrete symmetry case, vortex-like patterns are a direct consequence of the DMI.
@@ II.1 SC-tiling Intrinsic Scales
Figure 1(c) shows the so-called specific heat ($C$, see Appendix A) of the SC structure. Along with the plateau that reflects the 2D nature of the lattice, as $C=\frac{d}{2}$ for scale-invariant architectures [42], $C$ presents two possible small-scale modes or ultraviolet cut-offs (in the jargon of the Renormalization Group), $\Lambda_{1}$ and $\Lambda_{2}$, independent of the system size, plus a third peak at larger values of $\tau$ reflecting the whole lattice scale. Compared to a regular 2D lattice that only presents two peaks, here, the extra peak plays, as discussed below, a key role in determining, altering, and disrupting the phases and phase transitions of the system. The presence of multiple characteristic topological scales ($\Lambda_{1}$ and $\Lambda_{2}$), each associated with a different type of interaction (Heisenberg vs. DM-like), leads to frustration: if the system cannot simultaneously minimize both energy terms globally. As a consequence, domains settle into configurations that are locally stable but globally incompatible – a hallmark of frustrated systems. This manifests in metastability, slow dynamics, and history-dependent responses (see below).
Figure 2: Topological symmetry-breaking. (a) Temperature phase diagram of Eq.(1) computed using direct simulations on a SC lattice with $N=1024$ nodes ($J=1$). Polarized states are computationally detected with the polarization order parameter, $P$, while vortex phases are detected through the vortex order parameter $v$. The metastable phase corresponds to the broad region of divergent polarization susceptibility $\chi_{P}=N\sigma^{2}_{P}$. (b)-(c) Polarization ($P$) and vortex ($v$) order parameters versus temperature for different lattice sizes (see legend, $N=4L^{2}$) for (b) $J/D=1.4$ and (c) $J/D=0.8$. Inset: Rescaled temporal variance $\chi_{P}=N\sigma^{2}_{P}$ versus temperature, $T$ for different system sizes. The metastable phase also presents characteristic temporal oscillations in the global system polarization (see Appendix B). (d) Top: Far-field imaging of a SC structure at TT= $T_{C}$-2K. Bottom: Computational 2D Fourier transform of an ordered vortex phase (J/D=0.8,T=0.4,N=1024)J/D=0.8,~T=0.4,~N=1024). (e) Temperature phase diagram of the model versus electric field $E$ for $J/D=0.8$ and $N=1024$ nodes. The breakup of the vortex phase translates into the emergence of a metastable phase where the ferroelectric and vortex domains naturally coexist. (f)-(g) Polarization ($P$) and vortex ($v$) order parameters as a function of the electric field for different lattice sizes (as in (b) and (c)) for (f) $T=0.4$ and (g) $T=0.7$. The peak in the system susceptibility is indicated here as a vertical dashed line. Note how the critical field at low temperatures splits into two different phase transitions when T increases, with naturally coexisting vortices and ferroelectric domains. All curves have been averaged over $10^{2}-10^{3}$ independent realizations.
@@ II.2 Phase Diagram
Figure 2(a) shows the temperature phase diagram of the system as a function of the ratio between the dipole-dipole alignment interaction, $J$, and the DMI, $D$. Results are achieved through extensive Monte Carlo simulations of the TBM, running Eq. (1) on top of the SC-tiling. In order to reach zero local charge density, the constraint $\nabla\cdot\mathbf{p}_{\Lambda_{1}}=0$ is imposed as the energy is minimized (see Fig. 1(b)), centering it on each ”plaquette” (the elementary protected cell $\Lambda_{1}$) to avoid nonphysical situations (see SM [39] for case examples of microscopic configurations, resembling the ice rule in spin-ice systems [43]). Inspection of the phase diagram reveals two stable collective dynamical regimes: vortices and ferroelectric phases. In particular, the microscopic scale $\Lambda_{1}$ now supports the existence of an ordered domain mosaic of 2D protected vortex cores. Different types of bifurcation lines separate the phases of the system. In particular, for low temperatures, increasing the ratio $J/D$, a first-order phase transition separates the vortex phase from the ferroelectric one. Conversely, a second-order phase transition separates the vortex phase from the paraelectric one. Interestingly, a new phase emerges between the ferroelectric and paraelectric phases: a metastable, or frustrated, phase where transient vortices prevent the system from reaching a global order as in the ferroelectric or vortex phases, sequentially changing the full system orientation (see Appendix B).
@@
Figures 2(b) and (c) show order parameters as a function of the temperature for the polarization, $P$, and vortex, $v$, for two different values of the ratio $J/D$. Note that $P$ and $v$ correctly detect the stable ordered phases. Insets show the anomalous temporal fluctuations $\sigma^{2}_{P}$ and $\sigma^{2}_{v}$, which exhibit a pronounced peak located at the ($N$-dependent) transition point between the different phases. In particular, the broad divergence of $\sigma^{2}_{P}$ is a hallmark of frustrated phases in statistical mechanics supporting the picture of broad critical-like regions [44, 45], and give us a marker for the emergence of a new intermediate phase between the ferroelectric and the paraelectric ones (see also Appendix B).
@@
Figure 3: Direct imaging of SC ferroelectric clusters. Binarized crossed-polarizer transmission microscopy images for (a) $T_{C}$-4.5K and increasing electric field (0 kV/cm, 2.86 kV/cm, and 2.91 kV/cm, respectively, using $\theta=25$). Light transmission shows an abrupt transition at $E_{C}\approx$2.9 kV/cm. (b) Various conditions of temperature and bias electric field to enhance the critical point for the first two temperatures and the metastable region for $T_{C}$-2.5K (see images, with $\theta=37$, 3737, and $35$, respectively). (c-e) Light intensity versus applied electric field ($E$) for three different temperatures and different threshold values ($\theta$, see legend): (c) $T_{C}-4.5K$. An abrupt phase transition is present at a field $E_{C}\approx 2.9$ kV/cm for a wide range of $\theta$ values. (d) $T_{C}-3.5K$. The abrupt transition changes its nature to a continuous one with $E_{C}\approx$ 2.75 kV/cm and (e) $T_{C}-2.5K$. The system exhibits two different phase transitions at different critical fields, $E_{C}\approx$2.5 kV/cm and $E_{C}\approx$ 2.86 kV/cm (see also Appendix E).
@@
Findings are compared to experiments in SCs observed in KTN:Li crystals (see also Appendices C and D). This material, which led to the first observation of wave replica-symmetry-breaking [46], has been developed by several groups [47, 48, 49, 50] as a lead-free optically transparent alternative to other commonly used perovskites [51]. Moreover, compared to commercially available crystals, such as BTO [37] and SBN [52], it has tunable properties with a room-temperature ferroelectric transition [53, 54, 8]. Fourier computational analysis of the vortex phase (see also SM [39]) is found to match observed optical far-field results, as shown in Fig.2(d), in agreement with the previous SC models based on an interlocked regular structure of spontaneous polarization [6, 55, 54]. The TBM model then provides a feasible explanation to why the SC phase appears only in some specific materials, as the internal structure must be fine-tuned to present a specific $J/D$ ratio.